Jackie is twice as old as John. Ten years ago, Jackie was three times as old as John was ten years ago. How old is each one new? Is there a certain format to use when doing these type of word problems? I get lost on this. Thank you for any help you provide.

Answer by ankor@dixie-net.com(22740)

Jackie is twice as old as John.
Ten years ago, Jackie was three times as old as John was ten years ago.
How old is each one now?
:
Here is one method, that's pretty easy
:
Let x = Jackie's present age
Let y = John's present age
:
Write an equation for each statement:
:
"Jackie is twice as old as John."
x = 2y
:
Ten years ago, Jackie was three times as old as John was ten years ago.
x - 10 = 3(y - 10)
x - 10 = 3y - 30
x = 3y - 30 + 10
x = 3y - 20
:
From the 1st equation, replace x with 2y
2y = 3y - 20
+20 = 3y - 2y
y = 20 yrs is John's present age
then
x = 2(20)
x = 40 is Jackie's present age
:
:
Check solutions in the statement:
"Ten years ago, Jackie was three times as old as John was ten years ago."
40 - 10 = 3(20 - 10)
30 = 3(10), confirms our solutions